# Empirical Evaluation of Generalizable Automated Program Repair with Large Language Models

### v1

@@ IV-B1 Granularity of Fault Localization
We compare the base prompt – which includes (only) the faulty method and thus simulates perfect method-level FL – with the Line-level Localization (LL) prompt, which includes a hint about the particular code lines to fix. Table IV displays results for both prompts across the most promising models from RQ1. We see that line-level localization – while improving results – appears to add less value compared to test errors for (almost) all languages-model combinations. Also, we observe that the improvements with line-level localization appear to be inconsistent: Particularly, in the PHP benchmark, the accuracy deteriorates for out of LLMs when introducing line-level localization. This suggests that future research on localization in APR should evaluate across multiple programming languages.
IV-B2 Automated Fault Localization
@@
Assuming perfect (method- or line-level) localization may be a questionable assumption in practice, where the scope of bugs may have to be narrowed down within large software projects. To study the relationship between fault localization and neural patch generation in a realistic setting, we use FLACOCO [40], a recent spectrum-based localization tool, to identify the three most suspicious functions as potential locations for the fix. We then prompted the model using these candidate functions. As the tool currently supports only Java, this experiment was conducted on Defects4J (using the subselection of samples). The results are presented in Table VI. The first two columns show the plausible@1 results using perfect fault localization. In comparison, plausible@1 using automated fault localization in the third column shows a significant drop. The performance drop ranges from % for Gemini 2.0 Flash to % for Claude 3.7 Sonnet compared to the base prompt. This can be attributed to two main factors. First, the accuracy of the automated fault localization tool is lower than that of perfect fault localization. In our setup, the actual fix location appeared among the top-3 candidates for only out of faults. For the remaining cases, the prompt included only irrelevant code locations. We discovered only one such case for which two LLMs produced plausible patches, and even these could be attributed to test overfitting. Also, even when the correct function was identified, we observed a degradation in performance: on average, plausible@k decreased by %, and in % of those cases, no plausible patch could be generated – despite the base prompt yielding plausible solutions within the 15 generations per model. Since automated line-level localization can be expected to be even harder (and due to the mixed results with (perfect) line-level localization in Table IV), we did not explore automated line-level localization further.
@@
Overall, our results for RQ2 suggest that practical, automated localization has a significant impact on APR results. Future APR research might consider including practice-oriented evaluations with localization tools, or even put an even stronger focus on integrated approaches of repair and localization, such as [59].
@@
| Claude 3.7 Sonnet | 19.02% | 27.94% | 2.82% | |
| Gemini 1.5 Pro | 19.44% | 20.12% | 8.25% | |
| Gemini 2.0 Flash | 10.27% | 14.53% | 2.44% | |
| o3-mini | 15.10% | 23.11% | 3.64% | |
| DeepSeek R1 (dist.) | 17.16% | 18.83% | 0.31% | |
| Qwen 2.5 Coder | 10.11% | 12.75% | 2.79% | |
IV-C RQ3: Patch Complexity
We also test how repair accuracy varies for different levels of patch complexity. To do so, Table VII reports plausibility@1/5 for all benchmarks (using the ”test” prompt) separately for single-line, single-hunk, and multi-hunk issues. While we expect complexity to increase with these categories significantly, the expected performance deterioration was not as high as we would have expected. We observed a drop from (single-line) over (single-hunk) to (multi-hunk), i.e., LLMs cope still relatively well with increasing patch complexity. We even observe of performance figures to improve from single-line to single-hunk (particularly for Python), and of from single-hunk to multi-hunk (particularly for Java). Note that the experiment focused on single-function cases. However, these promising results suggest to even investigate multi-hunk, multi-function APR further.
@@ IV-C RQ3: Patch Complexity
| p@ | p@ | p@ | p@ | p@ | p@ | |||||
@@
IV-D RQ4: open vs closed Models
Unsurprisingly, the accuracy in most experiments is significantly higher on average for closed models, which are likely trained on larger and higher-quality datasets, are typically larger in parameter count, and may have been subject to more extensive fine-tuning on failure cases compared to open models. To get an impression of how open models and closed models progress over time, we plot the quality of program repair against the models’ release dates. As a quality measure, we use the plausible@5 rate (see Table II), averaged over all datasets. For OpenAI’s models, which undergo continuous development, we use the exact date of the version used in our experiments (for GPT-4o: Nov 11 2024, for o3-mini: Jan 31 2025).
@@ IV-D RQ4: open vs closed Models
V Discussion & Threats to Validity
@@
Recent research in APR [15, 16, 17] focused on building agentic workflows, which is not part of our empirical exploration. Instead, we focused on a more fundamental evaluation of LLMs for the under-explored area of cross-language repair capabilities. Compared to agentic workflows, we focus on single prompt setups using different prompt ingredients. Nevertheless, our insights will help other researchers and engineers select LLMs to design and build new agentic workflows that can handle cross-language repair scenarios.
@@
Sallou et al. [50] discussed several evaluation aspects when using LLMs in software engineering research, including the problem of output variability, data leakage, and closed-source models. To address the output variability of LLMs, we generated not only one but 15 patches (see details in Section III-D). To evaluate the generated patches, we used the established metric [52] representing an estimator for the model’s effectiveness. Assessing only the plausibility (i.e., passing test cases) and not the correctness (i.e., semantical equivalence with a ground truth like a developer patch) represents a threat to construct validity. However, manually assessing the actual correctness for all subjects is not feasible. The data leakage issue is more challenging to address because we cannot control the training data of state-of-the-art LLMs. Related works, e.g., RepairBench [42] used a dataset with more recent data, whose time period is now also included in the cut-off dates of the latest models. At this point, we have to assume that some of the benchmark data could be included in the training data of some of the LLMs we used in our experiments, which represents a threat to internal validity. In fact, Zhou et al. [60] recently investigated 83 SE benchmarks regarding their potential data leakage. According to their analysis Defects4J (in its 2.0 version) has a leaked ratio of 0.41% and BugsInPy 11.0%. Both numbers are relatively low. Interestingly, although BugsInPy obviously has some data leakage issues, our experiments showed it as the most challenging benchmark for the tested LLMs. As part of our empirical evaluation, we support a comparative analysis of models by selecting a wide range of open-source/weight and closed models. Our broad selection of models and benchmarks reduces the threat to external validity; however, we cannot claim generality beyond our experiments.
@@
As described in Section III-E, we used zero-shot prompts with different repair ingredients in our experiments. Those prompts were designed based on and in consistency with prior work [28, 22, 38, 42], and have been tested in preliminary experiments. Additional optimization and tuning of the prompts could lead to better repair performances on our benchmarks; however, we did intentionally not perform further optimizations to prevent introducing model-specific biases and further benchmark overfitting.
@@
We noticed that, e.g., due to its reasoning strategy, the DeepSeek R1 (dist.) model produces more iterative output, requiring significantly more time to conclude with the final repair than the other models. Other studies [35] showed that developers could accept a timeout of up to one hour for integrating APR technology in their workflows. This timeout is still much higher than LLMs would require; in our experiments spanning between seconds and several minutes. However, it is currently unknown what an acceptable timeout for LLM-based APR would be so that, e.g., it can effectively integrate into agentic workflows. Different resource demands across the models, in particular with respect to timing, can lead to unfair comparison representing a threat to internal validity. Our study methodology did not limit the time for model inference or patch validation because we wanted to focus on the actual repair capabilities, and we leave a more rigorous analysis regarding repair time for future work.
VI Conclusion
@@
VII Data Availability
The data supporting this study’s findings are available from the authors upon reasonable request.
VIII Disclaimer

### v2

@@ 4.2.1. Granularity of Fault Localization
We compare the base prompt – which includes (only) the faulty method and thus simulates perfect method-level FL – with the Line-level Localization (LL) prompt, which includes a hint about the particular code lines to fix. Table 4 displays results for both prompts across the most promising models from RQ1. We see that line-level localization – while improving results – appears to add less value compared to test errors for (almost) all languages-model combinations. Also, we observe that the improvements with line-level localization appear to be inconsistent: Particularly, in the PHP benchmark, the accuracy deteriorates for out of LLMs when introducing line-level localization. This suggests that future research on localization in APR should evaluate across multiple programming languages.
4.2.2. Automated Fault Localization
@@
Assuming perfect (method- or line-level) localization may be a questionable assumption in practice, where the scope of bugs may have to be narrowed down within large software projects (Liu et al., 2019a). To study the relationship between fault localization and neural patch generation in a realistic setting, we use FLACOCO (Silva et al., 2021), a recent spectrum-based localization tool, to identify the three most suspicious functions as potential locations for the fix. We then prompted the model using these candidate functions. As the tool currently supports only Java, this experiment was conducted on Defects4J (using the subselection of samples). The results are presented in Table 6. The first two columns show the pass@1 results using perfect fault localization. In comparison, pass@1 using automated fault localization in the third column shows a significant drop. The performance drop ranges from % for Gemini 2.0 Flash to % for Claude 3.7 Sonnet compared to the base prompt. This can be attributed to two main factors. First, the accuracy of the automated fault localization tool is lower than that of perfect fault localization. In our setup, the actual fix location appeared among the top-3 candidates for only out of faults. For the remaining cases, the prompt included only irrelevant code locations. We discovered only one such case for which two LLMs produced plausible patches, and even these could be attributed to test overfitting. Also, even when the correct function was identified, we observed a degradation in performance: on average, pass@k decreased by %, and in % of those cases, no plausible patch could be generated – despite the base prompt yielding plausible solutions within the 15 generations per model. Since automated line-level localization can be expected to be even harder (and due to the mixed results with (perfect) line-level localization in Table 4), we did not explore automated line-level localization further.
@@
Overall, our results for RQ2 suggest that practical, automated localization has a significant impact on APR results. Future APR research might consider including practice-oriented evaluations with localization tools, or even put an even stronger focus on integrated approaches of repair and localization, such as (Shariffdeen et al., 2025).
@@
| Claude 3.7 Sonnet | 19.02% | 27.78% | 2.76% | |
| Gemini 1.5 Pro | 19.36% | 20.36% | 3.33% | |
| Gemini 2.0 Flash | 10.27% | 14.25% | 2.40% | |
| o3-mini | 15.02% | 22.81% | 3.64% | |
| DeepSeek R1 (dist.) | 17.16% | 18.00% | 0.31% | |
| Qwen 2.5 Coder | 7.41% | 11.81% | 1.61% | |
4.3. RQ3: Patch Complexity
We also test how repair accuracy varies for different levels of patch complexity. To do so, Table 7 reports pass@1 and pass@5 for all benchmarks (using the test prompt) separately for single-line, single-hunk, and multi-hunk issues. While we expect complexity to increase with these categories significantly, the expected performance deterioration was not as high as we would have expected. For , we observed a drop from (single-line) over (single-hunk) to (multi-hunk), i.e., LLMs cope still relatively well with increasing patch complexity. We even observe of performance figures to improve from single-line to single-hunk (particularly for Python), and of from single-hunk to multi-hunk (particularly for Java). Note that the experiment focused on single-function cases. However, these promising results suggest to even investigate multi-hunk, multi-function APR further.
@@ 4.3. RQ3: Patch Complexity
| pass@ | pass@ | pass@ | pass@ | pass@ | pass@ | |||||
@@
4.4. RQ4: open vs closed Models
Unsurprisingly, the accuracy in most experiments is significantly higher on average for closed models, which are likely trained on larger and higher-quality datasets, are typically larger in parameter count, and may have been subject to more extensive fine-tuning on failure cases compared to open models. To get an impression of how open models and closed models progress over time, we plot the quality of program repair against the models’ release dates. As a quality measure, we use the pass@5 rate (see Table 2), averaged over all datasets. For OpenAI’s models, which undergo continuous development, we use the exact date of the version used in our experiments (for GPT-4o: Nov 11 2024, for o3-mini: Jan 31 2025).
@@ 4.4. RQ4: open vs closed Models
5. Discussion & Threats to Validity
@@
Recent research in APR (Zhang et al., 2024b; Ruan et al., 2024; Yang et al., 2024) focused on building agentic workflows, which is not part of our empirical exploration. Instead, we focused on a more fundamental evaluation of LLMs for the under-explored area of cross-language repair capabilities. Compared to agentic workflows, we focus on single prompt setups using different prompt ingredients. Nevertheless, our insights will help other researchers and engineers select LLMs to design and build new agentic workflows that can handle cross-language repair scenarios.
@@
Sallou et al. (Sallou et al., 2024) discussed several evaluation aspects when using LLMs in software engineering research, including the problem of output variability, data leakage, and closed-source models. To address the output variability of LLMs, we generated not only one but 15 patches (see details in Section 3.4). To evaluate the generated patches, we used the established metric (Chen et al., 2021) representing an estimator for the model’s effectiveness. Assessing only the plausibility (i.e., passing test cases) and not the correctness (i.e., semantical equivalence with a ground truth like a developer patch) represents a threat to construct validity. However, manually assessing the actual correctness for all subjects is not feasible. The data leakage issue is more challenging to address because we cannot control the training data of state-of-the-art LLMs. Related works, e.g., RepairBench (Silva and Monperrus, 2024) used a dataset with more recent data, whose time period is now also included in the cut-off dates of the latest models. At this point, we have to assume that some of the benchmark data may have been included in the training corpora of some of the LLMs we used in our experiments, which represents a threat to internal validity. In fact, Zhou et al. (Zhou et al., 2025) recently investigated 83 SE benchmarks regarding their potential data leakage. According to their analysis Defects4J (in its 2.0 version) has a leaked ratio of 0.41% and BugsInPy 11.0%. Both numbers are relatively low. Interestingly, although BugsInPy obviously has some data leakage issues, our experiments showed it as the most challenging benchmark for the tested LLMs. Ramos et al. (Ramos et al., 2025) further studied the impact of benchmark leakage across several open-source models and found that, although some models show substancial evidence of memorization on widely used benchmarks such as Defects4J, newer models, trained on larger datasets such as Llama 3.1 show only limited signs of leakage. As part of our empirical evaluation, we support a comparative analysis of models by selecting a wide range of open-source/weight and closed models. Our broad selection of models and benchmarks reduces the threat to external validity; however, we cannot claim generality beyond our experiments.
@@
As described in Section 3.5, we used zero-shot prompts with different repair ingredients in our experiments. Those prompts were designed based on and in consistency with prior work (Parasaram et al., 2024; Silva et al., 2024; Xia et al., 2023; Silva and Monperrus, 2024), and have been tested in preliminary experiments. Additional optimization and tuning of the prompts could lead to better repair performances on our benchmarks; however, we did intentionally not perform further optimizations to prevent introducing model-specific biases and further benchmark overfitting.
@@
We noticed that, e.g., due to its reasoning strategy, the DeepSeek R1 (dist.) model produces more iterative output, requiring significantly more time to conclude with the final repair than the other models. Other studies (Noller et al., 2022) showed that developers could accept a timeout of up to one hour for integrating APR technology in their workflows. This timeout is still much higher than LLMs would require; in our experiments spanning between seconds and several minutes. However, it is currently unknown what an acceptable timeout for LLM-based APR would be so that, e.g., it can effectively integrate into agentic workflows. Different resource demands across the models, in particular with respect to timing, can lead to unfair comparison representing a threat to internal validity. Our study methodology did not limit the time for model inference or patch validation because we wanted to focus on the actual repair capabilities, and we leave a more rigorous analysis regarding repair time for future work.
6. Conclusion
@@
7. Data Availability
The data that support the findings of this study, including all our results and the generated patches, are openly available in our supplemental material: https://figshare.com/s/947fd7030f10a67a1c9f. Upon acceptance, we will update the artifact with a reproduction package including the scripts for prompting the LLMs as well for the data analysis, and make the material accessible via a DOI.
8. Disclaimer